OpenStudy (anonymous):

Zebediah is an American traveling with a tour group in Eastern Europe. During a stop in Romania, he purchases a souvenir that is priced at 177 Romanian leu using his credit card. If the exchange rate that day is USD to RON = 2.89, which is the best estimate of the charge Zebediah will later find on his credit card statement?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

A USD is greater than RON BY 2.89: \[1*USD=2.89*1*RON\]RESULTS THAT: \[1*RON=\frac{ 1*USD }{ 2.89 }\] \[177*RON= \frac{ 177*USD }{ 2.89 }=61.24\]

OpenStudy (anonymous):

ur answer came out to be 61.24?

OpenStudy (anonymous):

that's not the answer

OpenStudy (anonymous):

then which is the answer